Solve the equation
40e^(-0.296x) = 100

Answer by lwsshak3(11628)

Solve the equation
40e^(-0.296x) = 100
e^(-0.296x) = 100/40=2.5
take log of both sides
(-0.296x)lne=ln(2.5)
-0.296x=ln(2.5)
x=ln(2.5)/-0.296
x≈-3.0955